# Experimental Study of Bremsstrahlung Gamma Ray Emission and Short-Range Correlations in 124Sn+124Sn Collisions at 25 MeV/u

### v1

@@ Abstract
Short-range correlation (SRC) in nuclei refers to the nucleons forming temporally correlated pairs in close proximity, giving rise to the high momentum of the nucleons beyond the Fermi surface. It has been reported that bremsstrahlung $\gamma$ production from neutron-proton process in heavy-ion reactions provides a potential probe of the existence of SRC in nuclei. In this paper, we present in detail the precision measurement of bremsstrahlung $\gamma$-rays in $\rm{}^{124}Sn$$\rm{}^{124}Sn$ reactions at 25 MeV/u using the Compact Spectrometer for Heavy IoN Experiment (CSHINE). The experimental setup, detector calibration, trigger scheme and data analysis procedures as well as the model comparison are presented in detail. Background contributions are subtracted using two methods to ensure robustness. By comparing the experimental $\gamma$ spectrum with the Isospin-dependent Boltzmann-Uehling-Uhlenbeck simulations, the high momentum tail (HMT) fraction of $R_{\rm HMT}=(20\pm 3)\%$ is derived in 124Sn nuclei. This work presents the detailed experimental measurement and analysis framework for the precise determination of the HMT fraction via bremsstrahlung $\gamma$-ray emission, demonstrating a new paradigm to study nucleon SRCs in nuclei using low-energy heavy-ion collisions.
@@ I Introduction
Short-range correlation (SRC) [1, 2, 3, 4] of nucleons is a temporal fluctuation that nucleons form temporally correlated pairs in close proximity in atomic nuclei. SRC plays an essential role in understanding the nuclear structure as well as to shed insight to the properties of dense nucleonic matter. Most traditional nuclear models are based on mean-field theory, in which nucleons are confined below the Fermi momentum and quark-level interactions are largely neglected. However, the EMC effect [5] indicates that deviations in nucleon momentum distributions from shell model predictions stem from the underlying quark structure of nucleons [6, 7, 8, 9, 10]. This is exemplified by the fact that nuclear parton distribution functions (PDFs) in heavy nuclei deviate markedly from the naive sum of those in free nucleons [11, 12, 13, 14, 15, 16], underscoring the importance of understanding the momentum distributions of quarks and gluons within bound nucleons and nuclei. SRCs, a ubiquitous feature of all nuclei, originate from fluctuations in the nuclear ground state and predominantly involve neutron-proton ($np$) pairs [17, 3, 18, 19, 20, 21, 22], driven by the strong tensor component of the nucleon-nucleon interaction at sub-Fermi-range distances [23, 24, 25]. Notably, $np$ pairs often display significant spatial overlap, which may favor isosinglet, spin-singlet $ud$ diquark configurations at the quark level [22]. These pairs also exhibit large relative momenta, contributing to the emergence of a high momentum tail (HMT) in the nucleon momentum distribution beyond the Fermi surface [19, 26]. Since SRCs reflect dynamics beyond the reach of mean-field approximations, incorporating both short-range nuclear forces and potential quark-level effects [27], they have become a focal point of contemporary theoretical and experimental research in nuclear physics.
@@
Hadronic reactions also serve as one of the primary approaches in the SRCs experimental studies. In exclusive scattering experiments involving large momentum transfer proton-nucleus collisions [38, 17], the momentum of the struck proton can be reconstructed while simultaneously measuring the momentum of the correlated neutron. Under high momentum transfer conditions, the struck fast-bound proton can be described using the instantaneous approximation, which significantly enhances the resolution of nuclear structure. Moreover, high-energy inverse kinematics scattering [39], where an incoming ion beam collides with a proton target, has emerged as a novel technique to investigate SRCs. By detecting the knocked-out protons and residual nuclear fragments, this method effectively suppresses complications from initial- and final-state interactions between hadrons [40, 41], thereby improving the reliability of reconstructed particle distributions.
@@
Heavy-ion collisions (HICs) provide the only means of producing and studying extreme states of nuclear matter in terrestrial laboratories. In the 1980s, high-energy photon emission observed in HICs emerged as a puzzling phenomenon [42, 43, 44], drawing growing interest due to its unclear origin. Subsequent experimental measurements have shown that a significant fraction of these $\gamma$-rays originate from bremsstrahlung processes induced by $np$ collisions in the early-stage of HICs [45, 46, 47, 48, 49, 50, 51, 52]. Recent advances in the study of SRCs have substantially deepened our understanding of nuclear structure. Theoretical studies indicate that high-energy bremsstrahlung $\gamma$-ray production in HICs [53] is sensitive to the high-momentum components of nucleon momentum distributions generated by SRCs [54], rendering hard photons in the $\gamma$-ray spectrum a promising probe of these correlations [53, 55]. Notably, photons interact only weakly with the nuclear medium and are largely unaffected by final-state interactions, making them a particularly clean and sensitive probe of short-range nuclear dynamics. Using the isospin-dependent Boltzmann-Uehling-Uhlenbeck (IBUU) transport model [56] with momentum-dependent interactions (MDI) [57], one can simulate $\gamma$-ray spectra from HICs by varying the fraction of high-momentum nucleons in the initial nuclear state. Comparing these simulations with experimental bremsstrahlung data enables the extraction of information on SRCs in nuclei.
@@
To explore the underlying dynamical mechanisms in intermediate-energy HICs, the Compact Spectrometer for Heavy Ion Experiment (CSHINE) [58, 59, 60] has been constructed, maintained, and continuously upgraded at the Radioactive Ion Beam Line in Lanzhou (RIBLL-1). Through multiple beam campaigns, CSHINE has developed advanced detection technologies and delivered a wide range of physics results in key areas including the equation of states of nuclear matter, isospin evolutions [61, 62, 63], short-range correlations in nuclei [64, 65, 66], neutron-neutron interactions [67], and the interplay between fission dynamics and isospin dynamics [68]. To date, CSHINE has completed five beam campaigns. As part of the detection system, a CsI(Tl)-based total absorption $\gamma$-ray spectrometer array (CSHINE-Gamma) [69] was first deployed during the fourth experiment in 2022, which involved 25 MeV/u 86Kr + 124Sn reactions. In this experiment, CSHINE-Gamma successfully detected and analyzed $\gamma$-ray spectra emitted from HICs. The data revealed an nonzero HMT fraction, providing experimental indication for the presence of SRC-induced HMT components in nuclei at about 90% confidence level [66, 65]. However, this early study faced limitations due to the restricted energy coverage of CSHINE-Gamma and a narrow event reconstruction region confined to the central scintillator. These constraints hindered the ability to quantify the HMT fraction with high statistical confidence. Moreover, due to the use of an asymmetric reaction system, it was not possible to determine the SRC fraction for a specific nucleus.
@@
In 2024, the fifth CSHINE experiment, involving a 25 MeV/u $\rm{}^{124}Sn$$\rm{}^{124}Sn$  beam, introduced significant upgrades of CSHINE-Gamma specifically aimed at enhancing high-energy $\gamma$-ray detection. These improvements included an extended energy range, expanded event reconstruction regions, and overall enhanced detection efficiency of high-energy $\gamma$-rays. As a result, the statistics for high-energy bremsstrahlung $\gamma$-rays were significantly increased. Leveraging the unprecedented statistical precision and improved measurement accuracy of this experiment, the HMT fraction in the nucleon momentum distribution of 124Sn nuclei has been determined quantitatively with improved precision [64]. This result marks the first high-precision extraction of the SRC fraction in low-energy HICs, achieved using bremsstrahlung $\gamma$-rays as a clean and unambiguous probe.
In this paper, we present the analysis and results of the 25 MeV/u $\rm{}^{124}Sn$$\rm{}^{124}Sn$  experiment, providing in detail the technical underpingnings of a submitted letter [64]. Specifically, we describe the detector setup, calibration and data analysis procedure. The final bremsstrahlung $\gamma$ spectra were derived using different methods, all of which were compared with the transport model simulations, yielding a consistent conclusion on the fraction of SRC in 124Sn nuclei. This paper is organized as following: Section II outlines the experimental setup, including the CSHINE detectors and $\gamma$ hodoscope calibration. Section III details the data analysis procedure. Section IV describes the IBUU-MDI model and parameter tests theoretically. Section V presents results and discussion, focusing on data uncertainties and HMT fraction analysis of measured $\gamma$ spectra. Section VI reconstructs the original $\gamma$ spectrum using Richardson-Lucy deblurring method and compares it with IBUU-MDI predictions directly. Section VII concludes with a summary and outlook.
@@ II.1 Detector Configuration of CSHINE
Figure 2: (Color Online) Trigger scheme of the $\rm{}^{124}Sn$$\rm{}^{124}Sn$  beam experiment.
@@
The neutrons are measured by a plastic scintillator array consisting of $4\times 5$ units, covering the laboratory polar angle $17^{\circ}<\theta_{\rm lab}<53^{\circ}$ in partial azimuth. Each unit is a block of plastic scintillator with the size of $\rm 15\times 15\times 15~cm^{3}$ readout by photomultiplier tube (PMT). The distance of the neutron array to the target is 200 cm. The neutron energies are measured by the time of flight (TOF) method. The start timing of TOF is provided by 4 $\rm BaF_{2}$ fast scintillators surrounding the target fired by the reaction $\gamma$-rays. For the performance of the neutron array, one can refer to [71]. Besides the neutron wall, a liquid scintillator neutron detector, with the sensitive volume of $\rm\phi 34.8~cm\times 102~cm$, was installed at 513 cm to the target at $\theta_{\rm lab}=60^{\circ}$. Two 8-inch PMTs are attached to the left and right ends of the sensitive volume.
@@
The bremsstrahlung $\gamma$-rays of this analysis are measured by a CsI(Tl) hodoscope (CSHINE-Gamma), containing 15 units in a $4\times 4$ configuration (with one corner being vacant). Each CsI(Tl) unit has the dimension of $\rm 7\times 7\times 25~cm^{3}$ and is coupled to a PMT of Hamamatsu R2631 for signal readout. The radiation length of CsI(Tl) is $X_{0}=8.39\rm~g/cm^{2}$ and the Moliere radius is $R_{\rm M}=3.531~\rm cm$. The energy resolution of the units is about $3.6\%$ for 1 MeV $\gamma$-ray and about 2% for $\gamma$-ray energy beyond 10 MeV. In order to suppress the cosmic-ray muons, 3 thick plastic scintillators ($5\times 30\times 30\rm~cm^{3}$) were mounted surrounding the CsI(Tl) hodoscope on top, left and right sides. The linearity of the CsI(Tl) response to high-energy $\gamma$-rays was tested after the experiment using the quasi monochromatic $\gamma$ beam at the Shanghai Laser Electron Gamma Source (SLEGS) [72]. The performance of the $\gamma$ hodoscope can be found in [69].

### v2

@@ Abstract
Short-range correlation (SRC) in nuclei refers to nucleons forming temporally correlated pairs in close proximity, giving rise to the high momentum of the nucleons beyond the Fermi surface. It has been reported that bremsstrahlung $\gamma$ production from neutron-proton process in heavy-ion reactions provides a potential probe to the SRC abundance in nuclei. In this paper, we present in detail the precision measurement of bremsstrahlung $\gamma$-rays in $\rm{}^{124}Sn$$\rm{}^{124}Sn$ reactions at 25 MeV/u using the Compact Spectrometer for Heavy IoN Experiment (CSHINE). A comprehensive experimental and analysis framework is established to ensure the reliability and robustness of the extracted results. Background contributions are evaluated and subtracted using independent methods, and the consistency of the analysis is systematically validated. By comparing the experimental $\gamma$ spectrum with the Isospin-dependent Boltzmann-Uehling-Uhlenbeck simulations, the high momentum tail (HMT) fraction of $R_{\rm HMT}=(20\pm 3)\%$ is derived in 124Sn nuclei. This work provides a detailed and validated experimental framework for extracting SRC information from bremsstrahlung $\gamma$-ray emission and demonstrates the feasibility of studying nucleon SRCs with high precision in low-energy heavy-ion collisions.
@@ I Introduction
Short-range correlation (SRC) [25, 12, 76, 2] of nucleons is a temporal fluctuation in which nucleons form temporally correlated pairs in close proximity in atomic nuclei. SRC plays an essential role in understanding the nuclear structure as well as to shed insight to the properties of dense nucleonic matter. Most traditional nuclear models are based on mean-field theory, in which nucleons are confined below the Fermi momentum and quark-level interactions are largely neglected. However, the EMC effect [4] indicates that deviations in nucleon momentum distributions from shell model predictions stem from the underlying quark structure of nucleons [27, 40, 57, 17, 51]. This is exemplified by the fact that nuclear parton distribution functions (PDFs) in heavy nuclei deviate markedly from the naive sum of those in free nucleons [44, 22, 32, 1, 71, 67], underscoring the importance of understanding the momentum distributions of quarks and gluons within bound nucleons and nuclei. SRCs, a ubiquitous feature of all nuclei, originate from fluctuations in the nuclear ground state and predominantly involve neutron-proton ($np$) pairs [62, 76, 42, 34, 19, 43, 89], driven by the strong tensor component of the nucleon-nucleon interaction at sub-Fermi-range distances [69, 90, 91]. Notably, $np$ pairs often display significant spatial overlap, which may favor isosinglet, spin-singlet $ud$ diquark configurations at the quark level [89]. These pairs also exhibit large relative momenta, contributing to the emergence of a high-momentum tail (HMT) in the nucleon momentum distribution beyond the Fermi surface [34, 18]. Since SRCs reflect dynamics beyond the reach of mean-field approximations, incorporating both short-range nuclear forces and potential quark-level effects [16], they have become a focal point of contemporary theoretical and experimental research in nuclear physics.
@@
Hadronic reactions also serve as one of the primary approaches for experimental studies of SRCs. In exclusive scattering experiments involving large momentum transfer proton-nucleus collisions [77, 62], the momentum of the struck proton can be reconstructed while the momentum of the correlated neutron is measured simultaneously. Under high momentum transfer conditions, the struck fast-bound proton can be described using the instantaneous approximation, which significantly enhances the resolution of the nuclear structure. Moreover, high-energy inverse kinematics scattering [61], where an incoming ion beam collides with a proton target, has emerged as a novel technique to investigate SRCs. By detecting the knocked-out protons and residual nuclear fragments, this method effectively suppresses complications from initial- and final-state interactions between hadrons [80, 13], thereby improving the reliability of reconstructed particle distributions.
@@
Heavy-ion collisions (HICs) provide the only means of producing and studying extreme states of nuclear matter in terrestrial laboratories. In the 1980s, high-energy photon emission observed in HICs emerged as a puzzling phenomenon [56, 55, 75], drawing growing interest due to its unclear origin. Subsequent experimental measurements have shown that a significant fraction of these $\gamma$-rays originate from bremsstrahlung processes induced by $np$ collisions in the early stage of HICs [7, 6, 65, 41, 38, 81, 78, 53]. Recent advances in the study of SRCs have substantially deepened our understanding of nuclear structure. Theoretical studies indicate that high-energy bremsstrahlung $\gamma$-ray production in HICs [98] is sensitive to the high-momentum components of nucleon momentum distributions generated by SRCs [93], rendering hard photons in the $\gamma$-ray spectrum a promising probe of these correlations [98, 31]. Notably, compared to the hadron probes proposed [87, 37, 39], photons interact only weakly with the nuclear medium and are largely unaffected by final-state interactions, making them a particularly clean and sensitive probe of short-range nuclear dynamics. Using the isospin-dependent Boltzmann-Uehling-Uhlenbeck (IBUU) transport model [45] with momentum-dependent interactions (MDI) [15], one can simulate $\gamma$-ray spectra from HICs by varying the fraction of high-momentum nucleons in the initial nuclear state. Comparing these simulations with experimental bremsstrahlung data enables the extraction of information on SRC fractions in nuclei.
@@
To explore the underlying dynamical mechanisms in intermediate-energy HICs, the Compact Spectrometer for Heavy Ion Experiment (CSHINE) [28, 82, 88] has been constructed, maintained, and continuously upgraded at the Radioactive Ion Beam Line in Lanzhou (RIBLL-1). Through multiple beam campaigns, CSHINE has developed advanced detection technologies and delivered a wide range of physics results in key areas including the equation of states of nuclear matter, isospin evolutions [101, 84, 85], short-range correlations in nuclei [94, 95, 64], neutron-neutron interactions [73], and the interplay between fission dynamics and isospin dynamics [83]. To date, CSHINE has completed five beam campaigns. As part of the detection system, a CsI(Tl)-based total absorption $\gamma$-ray spectrometer array (CSHINE-Gamma) [63] was first deployed during the fourth experiment in 2022, which involved 25 MeV/u 86Kr + 124Sn reactions. In this experiment, CSHINE-Gamma successfully detected and analyzed $\gamma$-ray spectra emitted from HICs. The data revealed a non-zero HMT fraction, providing experimental indication of the presence of SRC-induced HMT components in nuclei at about 90% confidence level [64, 95]. However, this early study faced limitations due to the restricted energy coverage of CSHINE-Gamma and a narrow event reconstruction region confined to the central scintillator. These constraints hindered the ability to quantify the HMT fraction with high statistical confidence. Moreover, due to the use of an asymmetric reaction system, it was not possible to determine the SRC fraction for a specific nucleus.
@@
In 2024, the fifth CSHINE experiment, involving the 25 MeV/u $\rm{}^{124}Sn$$\rm{}^{124}Sn$  reactions, introduced significant upgrades to CSHINE-Gamma specifically improving the high-energy $\gamma$-ray detection. These improvements included an extended energy range, expanded event reconstruction regions, and overall enhanced detection efficiency of high-energy $\gamma$-rays. As a result, the statistics for high-energy bremsstrahlung $\gamma$-rays were significantly increased. Leveraging the unprecedented statistical precision and improved measurement accuracy of this experiment, the HMT fraction in the nucleon momentum distribution of 124Sn nuclei has been determined quantitatively with improved precision [94]. This result marks the first statistically high-precision extraction of the SRC fraction in low-energy HICs, achieved using bremsstrahlung $\gamma$-rays as a clean and sensitive probe.
In this paper, we present a comprehensive analysis of the SRC fraction extracted from the bremsstrahlung $\gamma$ spectrum in the 25 MeV/u $\rm{}^{124}Sn$$\rm{}^{124}Sn$  experiment. Compared to the short article reporting the primary physics outcome [94], the present work emphasizes a systematic validation of the analysis and a transparent documentation of the methodologies employed. Specifically, we describe the detector setup, calibration procedures, and data analysis workflow in detail, and assess the robustness of the extracted SRC fraction through multiple independent and complementary approaches. On the theoretical side, a complete formulation of the IBUU-MDI transport model is presented together with systematic studies of parameter dependencies, demonstrating that the HMT fraction plays a dominant role in shaping the high-energy bremsstrahlung $\gamma$ spectrum. The SRC fraction is independently extracted using different coincidence selections, which yield consistent results within uncertainties and are incorporated into the systematic error budget. In addition, the bremsstrahlung $\gamma$ spectrum is reconstructed using an independent unfolding approach, enabling a direct comparison with theoretical predictions without detector-response folding. These combined analyses establish the robustness and reliability of the extracted HMT fraction and extend significantly beyond the scope of Ref. [94].
This paper is organized as follows: Section II outlines the experimental setup. Section III describes the IBUU-MDI transport model and the corresponding theoretical studies. Section IV focuses on data uncertainties and the extraction of the HMT fraction from the measured $\gamma$ spectra, including validation using different coincidence selections. Section V presents an independent reconstruction of the bremsstrahlung $\gamma$ spectrum and its direct comparison with model predictions. Finally, Section VI concludes with a summary and outlook. The detailed descriptions of the trigger scheme, detector calibration, event reconstruction procedures, and background evaluation are provided in the Supplementary Material.
@@ II Experimental Setup
The neutrons are measured by a plastic scintillator array consisting of $4\times 5$ units, covering the laboratory polar angle $17^{\circ}<\theta_{\rm lab}<53^{\circ}$ in partial azimuth. Each unit is a block of plastic scintillator with the size of $\rm 15\times 15\times 15~cm^{3}$ read out by a photomultiplier tube (PMT). The distance of the neutron array to the target is 200 cm. The neutron energies are measured by the time of flight (TOF) method. The start timing of TOF is provided by 4 $\rm BaF_{2}$ fast scintillators surrounding the target fired by the reaction $\gamma$-rays. For the performance of the neutron array, one can refer to [74]. Besides the neutron wall, a liquid scintillator neutron detector, with the sensitive volume of $\rm\phi 34.8~cm\times 102~cm$, was installed at 513 cm to the target at $\theta_{\rm lab}=60^{\circ}$. Two 8-inch PMTs are attached to the left and right ends of the sensitive volume.
@@
The bremsstrahlung $\gamma$-rays of this analysis are measured by a CsI(Tl) hodoscope (CSHINE-Gamma), containing 15 units in a $4\times 4$ configuration (with one corner being vacant). Each CsI(Tl) unit has the dimension of $\rm 7\times 7\times 25~cm^{3}$ and is coupled to a PMT of Hamamatsu R2631 for signal readout. The radiation length of CsI(Tl) is $X_{0}=8.39\rm~g/cm^{2}$ and the Moliere radius is $r_{\rm M}=3.531~\rm cm$. The energy resolution of the units is about $3.6\%$ for 1 MeV $\gamma$-ray and about 2% for $\gamma$-ray energy beyond 10 MeV. In order to suppress the cosmic-ray muons, 3 thick plastic scintillators ($5\times 30\times 30\rm~cm^{3}$) were mounted surrounding the CsI(Tl) hodoscope on top, left and right sides. The linearity of the CsI(Tl) response to high-energy $\gamma$-rays was tested after the experiment using the quasi-monochromatic $\gamma$ beam at the Shanghai Laser Electron Gamma Source (SLEGS) [97]. The performance of the $\gamma$ hodoscope can be found in [63].